Circular No. 6186

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

PKS 1622–297 J. T. Bonnell and C. R. Shrader, Goddard Space Flight Center (GSFC)

and Universities Space Research Association; R. C. Hartman, GSFC; W. T. Vestrand and J. G. Stacy, Space Science Center, University of New Hampshire; and J. R. Webb, Florida International University, communicate: “We detected the quasar PKS 1622–297 with the International Ultraviolet Explorer (IUE) satellite using the Long Wavelength camera in low-resolution mode during its period of high-level activity in high- and low-energy γ-rays (IAUC 6185). The flux at 260 nm was ∼ 10^{−15} erg cm^{−2} s^{−1} A^{−1} on July 1. Subsequent IUE observations of this source on July 5, 6, and 8 suggest variability with the peak observed flux level occurring on July 8. Further IUE observations are planned to coincide with monitoring by the OSSE instrument onboard the Compton Gamma Ray Observatory (through July 24). Observations at other wavelengths continue to be very important.”

VARIABLE STAR IN SERPENS K. W. Hodapp, University of Hawaii, reports: “Observations show that

a young star located at α = 18^{h}27^{m}16^{s}.8, δ = +1^{o}14′15′′ (equinox 1950.0; uncertainty ± 2′′), deeply embedded in the Serpens molecular cloud, brightened by 4.1 mag in the K band (2.2 μm) in the past year. The following K magnitudes were obtained with the University of Hawaii 2.2-m telescope (+ QUIRC 1024×1024 infrared camera): 1994 Aug. 15–17 UT, 17.7 ± 0.3; 1995 July 12, 13.6 ± 0.3. (The position and magnitudes are relative to stars observed by Eiroa and Casali 1992, A.Ap. 262, 468). The star now appears associated with a nebula extending 10′′ to the west. Imaging at wavelengths of 2.06, 2.12, and 2.26 μm indicates that this nebula is primarily continuum radiation, probably scattered light. The star is invisible in the H band (1.6 μm), even though the nebula remains partly visible; both the star and the nebula are invisible in J (1.2 μm). The observed increase in flux from this very young star is probably a FU Ori-type outburst caused by an increase in mass accretion and a resulting increase in disk luminosity.”

GRO J1629–49 S. N. Zhang, B. A. Harmon, W. S. Paciesas, and G. J. Fishman report,

on behalf of the Compton Gamma Ray Observatory BATSE team, that this source (IAUC 6182) was not detectable by BATSE between June 29 and July 7. They set a 3σ upper limit of 50 mCrab between 20 and 100 keV.

1995 July 13 Daniel W. E. Green